| ID | 
Date | 
Author | 
Category | 
Subject | 
 
| 
  291  
 | 
Mon Apr  7 19:41:31 2025 | 
Oliver, Filip | General | Lassie screenshots |   |  
| 
  290  
 | 
Mon Apr  7 18:58:04 2025 | 
Shahab | DAQ | changed averaging in NPZ files |  new analysis settings for NPZ files:
[analysis]
nframes = 190
lframes = 65536
navg = 1
first file with this setting:
'RSA02-2025.04.07.18.57.17.454.tiq', 'RSA02-2025.04.07.18.57.02.328.tiq', 'RSA02-2025.04.07.18.58.17.958.tiq', 'RSA02-2025.04.07.18.57.40.143.tiq', 'RSA02-2025.04.07.18.58.25.520.tiq', 
'RSA02-2025.04.07.18.57.09.892.tiq', 'RSA02-2025.04.07.18.57.32.580.tiq', 'RSA02-2025.04.07.18.57.25.017.tiq', 'RSA02-2025.04.07.18.58.10.395.tiq', 'RSA02-2025.04.07.18.57.55.271.tiq', 
'RSA02-2025.04.07.18.56.54.765.tiq', 'RSA02-2025.04.07.18.58.02.833.tiq', 'RSA02-2025.04.07.18.57.47.707.tiq', 'RSA02-2025.04.07.18.56.47.202.tiq'  |  
| 
  289  
 | 
Mon Apr  7 18:52:15 2025 | 
Shahab, Oliver | DAQ | updated DAQ settings |  changed 410 resonance frequency back to plunger 8
245 MHz plunger stays at 350
center freq:
410 at 414.474 MHz
245 at 244.000 MHz
first good files in this setting are 18:51 onwards.
everything else remains same like last time.  |  
| 
  288  
 | 
Mon Apr  7 18:42:17 2025 | 
Oliver, Filip | General | Lassie screenshots |   |  
| 
  287  
 | 
Mon Apr  7 18:37:41 2025 | 
Shahab, Oliver | DAQ | updated DAQ settings |  changed 410 resonance frequency back to 212th harmonic
plunger 33
center freq:
410,600 MHz
first good files in this setting are 18:36 onwards.
everything else remains same like last time.  |  
| 
  286  
 | 
Mon Apr  7 18:17:41 2025 | 
Shahab, Oliver | DAQ | updated DAQ settings |  changed 410 resonance frequency back to 213th harmonic
plunger 20
center freq:
412.537 MHz
first good files in this setting are 18:16 onwards.
everything else remains same like last time.  |  
| 
  285  
 | 
Mon Apr  7 17:51:15 2025 | 
Shahab | General | HowTo change resonance frequency of 410 MHz Cavity |  step by step pictures:
sometimes the name of the motor connection is not available in the menu, it should be added manually.
Add MBOX:
put some name, like "410 MHz Schottky Cavity in ESR"
put the IP address: 140.181.139.194
press connect
change to local control
press initialize
go to motor drive
go to motor 1 and 2
always move same value for both motors.  |  
| 
  284  
 | 
Mon Apr  7 17:26:57 2025 | 
Shahab, Oliver | DAQ | New DAQ Settings |  new DAQ Settings for the RSAs:
414475588 (192Pb+82   at 214 harmonic)
244.03689
plunger settings: 410 MHz cavity--> 8
245MHz cavity --> 350
recording at 50 MSPS
2^16 points each frame -->  1ms per frame
Average set to: 5 frames --> total of ca. 6.5 ms per average frame
first good NPZ file from this settings on RSAs starts from 17:41  |  
| 
  283  
 | 
Mon Apr  7 17:25:00 2025 | 
Oliver, Filip | General | Lassie screenshots |   |  
| 
  282  
 | 
Mon Apr  7 17:18:45 2025 | 
Wolfram, Yuri, Sergey | General | Data from last 10 minutes are not useful (setting ESR) |   |  
| 
  281  
 | 
Mon Apr  7 17:07:33 2025 | 
Team PLEIADES | Detectors | Moving OUTSIDE PLEIADES detector slowly in (and out) |  17:08 Tried to move outside particle detector (GE01DD2AG) in - it is stuck again :-(
17:18 Sergey tried to move it, now it moves again. Moved to -100 mm.
17:30 Moved detector to outside parking position  |  
| 
  279  
 | 
Mon Apr  7 16:19:30 2025 | 
Oliver, Filip | General | Lassie screenshots |   |  
| 
  278  
 | 
Mon Apr  7 16:11:44 2025 | 
Ruijiu Chen, Yuri, Wolfram, | Analysis | PID of 192Pb+82 |   |  
| 
  277  
 | 
Mon Apr  7 16:07:42 2025 | 
Shahab | General | all config files |  here are all TOML config files of the automatic scripts.  |  
| 
  276  
 | 
Mon Apr  7 15:47:03 2025 | 
Shahab, Taka, Ruijiu | DAQ | change NAS and restart both acquisitions |  changed storage place to bank 1 of the NAS
restarted the NTCAP with new address
restarted Analyzers with new address
restarted the queqiao script
restarted looper scripts
new files on RSAs start from:
2025-04-07 15:23 onwards
new files on NTCAP have their corresponding time stamp on the directory  |  
| 
  275  
 | 
Mon Apr  7 15:23:01 2025 | 
Ragan, Wolfram, Sergey, Zsolt, Yuri, Rui-Jiu | General | ESR Setting change to 192Pb82+ |  Today, until 11 am, data were taken for the 188Hf setting. Since then, we are now tuning on the 192Pb82+.  |  
| 
  274  
 | 
Mon Apr  7 15:14:54 2025 | 
Oliver, Filip | General | lassie screenshots |   |  
| 
  273  
 | 
Mon Apr  7 14:05:40 2025 | 
Jelena, Guy | General | Lassie screenshots |   |  
| 
  272  
 | 
Mon Apr  7 13:18:27 2025 | 
Jelena, Guy | General | Tuning |  Beam was gone at 13:04 due to fragment tuning.  |  
| 
  271  
 | 
Mon Apr  7 12:29:16 2025 | 
Jelena, Guy | General | Lassie screenshots |   |  
 
 |